OpenStudy (anonymous):

The length of the base of an isosceles triangle is 4 inches less than the length of one of the two equal sides of the triangle if the perimeter is 32.

OpenStudy (anonymous):

length = 12

OpenStudy (anonymous):

let x be the length of the base y be the length of one of the 2 equal sides x=y-4 x+y+y=32 y-4+y+y=32 3y=36 y=12 x=8

OpenStudy (anonymous):

okie dokie... thanks guys :D